# A partition cover approach to tokenization

### v2

@@ 4.1 Challenges in designing an efficient algorithm for Problem 1
Informally speaking, submodular functions represent diminishing returns while supermodular functions represent increasing returns. In the context of the tokenization problem, the set $\mathbf{T}$ represents the candidate set of all possible tokens. Unfortunately for us, Problem 1 is neither submodular nor supermodular; see Table 1.
@@
[…]
@@
$20=f(\mathbf{X}\cup\{Z\})-f(\mathbf{X})>f(\mathbf{Y}\cup\{Z\})-f(\mathbf{Y})=-$ 2 $-1=f(\mathbf{X}\cup\{Z^{\prime}\})-f(\mathbf{X})<f(\mathbf{Y}^{\prime}\cup\{Z^{\prime}\})-f(\mathbf{Y}^{\prime})=0$
[…]
The second approach is to relax the original problem into one that is efficiently solvable and then prove approximation guarantees by relating between the relaxed solution to the original problem. For example, one can formulate the NP-hard vertex cover problem as an integer linear program, relax the integral constraints, solve the resulting linear program in polynomial time, and round the fractional solution back into an integral solution that corresponds to an actual vertex cover. This approach yields a 2-approximation algorithm for the vertex cover problem, e.g. see [WS11, Section 1.2]. Alas, real-world NLP problems involve large corpora which translates to problem instances with an extremely large number of variables and constraints which makes this relax-and-round paradigm practically infeasible.
4.2 An equivalent mixed integer program formulation
To design our algorithm GreedTok for the tokenization problem, we first show an equivalent formulation of Problem 1 in terms of a mixed integer program (MIP). The rationale behind this is two-fold. First, the MIP formulation offers a more straightforward and intuitive framework for defining our greedy algorithm. This clarity simplifies the process of understanding and implementing the algorithm within the given constraints. Second, the MIP formulation naturally lends itself to a relaxation to the well-established weighted maximum coverage problem (WMC), which is known to be submodular and has a corresponding greedy algorithm GreedWMC that provably achieves an $(1-1/e)$-approximation; see [Hoc96, Section 3.9]. Although we cannot formally establish approximation guarantees for GreedTok, its connection to WMC enables empirical comparisons with GreedWMC in Tok instances; see Section 5.3.
To begin, let us define $\textsc{cover}(W,\mathbf{S})$ as the maximum number of adjacent singletons in word $W$ that can be grouped into tokens from $\mathbf{S}$, with each character used at most once. For example, with $W=\text{scaredy}$ and $\mathbf{S}=\{\text{care},\text{edy}\}$, we have $\textsc{cover}(W,\mathbf{S})=3$ from concatenating 3 adjacent singleton pairs in “care”, constrained on the position of ‘e’ which can only be used once. Meanwhile, $\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})=4$ via scaredy. Notice that $|W|=\textsc{Partition}+\textsc{Cover}$. This lets us rewrite the minimization objective from Problem 1 as an equivalent maximization objective:
Henceforth, we use Tok to refer to both minimization and maximization versions of the tokenization problem.
[…]
To formulate Problem 1 as an MIP, our goal is to choose a subset $\mathbf{S}\subseteq\mathbf{T}$ of size $|\mathbf{S}|\leq k$ such that the following objective is maximized, encoding $\max\sum_{W\in\mathbf{W}}\textsc{count}(W)\cdot\textsc{cover}(W,\mathbf{S})$:
@@ 4.2 An equivalent mixed integer program formulation
We remark that the objective Eq. 1 can be re-expressed as $\max\sum_{T\in\mathbf{T}}\sum_{W\in\mathbf{W}}\sum_{i=1}^{|W|-1}c_{W}\cdot m^{W,T}_{i,i+1}$, making Eq. 4 redundant. However, this current formulation is useful for showing how to relax Tok to WMC later.
Now, let us interpret and explain the constraints. Eq. 2 models the constraint that we are choosing a subset of size $|\mathbf{S}|\leq k$. Eq. 3 models the constraint that we can only use $T\in\mathbf{T}$ to cover if it is chosen in $\mathbf{S}$. Eq. 4 models the constraint that if a cover happened between two adjacent singletons, then a relevant $T\in\mathbf{T}$ must have been chosen in $\mathbf{S}$. However, Eq. 5 models the constraint of only covering two adjacent singletons with a single relevant $T\in\mathbf{S}$. Eq. 6 models the constraint of covering the entire substring $T\in\mathbf{T}$, or leave it uncovered. Eq. 7 and Eq. 8 model the constraints preventing the chosen substring $T\in\mathbf{T}$ from sharing the cover with another partially overlapping $T$.
@@
Consider the word $W=\text{ababa}$ and the token $T=\text{aba}$ has $x_{T}=1$, i.e. $T\in\mathbf{S}\subseteq\mathbf{T}$. If we only use $T$ to cover singletons in $W$ with left-to-right priority, then the resultant tokenized form of $W$ is ababa. So, $m^{W}=(1,1,0,0)$, $m^{W,T}=(1,1,0,0)$, and $m^{W,T^{\prime}}=(0,0,0,0)$ for all $T^{\prime}\in\mathbf{T}\setminus\{T\}$. Observe that the 00 bits in $m^{W}$ precisely denote the partitioning positions within $W$. Furthermore, the constraints Eq. 5 and Eq. 6 ensure that $T$ is the only token that occupies the first two adjacent singletons, while constraints Eq. 7 and Eq. 8 prevent an invalid overlap of $T$ for the last two adjacent singletons. Now, suppose if we also have $T^{\prime}=\text{ba}$ with $x_{T^{\prime}}=1$. Using both $T$ and $T^{\prime}$ to tokenize $W$ results in ababa with $m^{W}=(1,1,0,1)$, $m^{W,T}=(1,1,0,0)$, $m^{W,T^{\prime}}=(0,0,0,1)$, and $m^{W,T^{\prime\prime}}=(0,0,0,0)$ for all $T^{\prime\prime}\in\mathbf{T}\setminus\{T,T^{\prime}\}$.
@@
4.2.1 Relation to the weighted maximum coverage problem
Like the vertex cover problem, the weighted maximum coverage problem (WMC) is NP-hard problem [Kar72, WS11, Hoc96]. Given a set of unique elements $\mathbf{L}=\{L_{1},\dots,L_{|\mathbf{L}|}\}$ and their corresponding weights $\mathcal{W}=\{w_{1},\dots,w_{|\mathbf{L}|}\}$, a collection of sets $\mathbf{U}=\{U_{1},\dots,U_{|\mathbf{U}|}\}$ where $U\in\mathbf{U}\subseteq\mathbf{L}$, and a number $k$, we want to find a subset $\mathbf{U}^{\prime}\subseteq\mathbf{U}$ such that $|\mathbf{U}^{\prime}|\leq k$ and the total weights of covered elements $\sum_{L_{i}\in\bigcup\mathbf{U}^{\prime}}w_{i}$ is maximized. Formulating WMC as a mixed integer program, we have the objective:
@@ 4.2.1 Relation to the weighted maximum coverage problem
Let us now interpret and explain the constraints. Eq. 10 limits the number of selected sets $\leq k$. Eq. 11 ensures that if an element is covered, at least one of the sets containing the element must be included in $\mathbf{U}^{\prime}$. To see that WMC is a relaxation of Tok, we first establish a mapping between the variables between Tok and WMC:
  • $m^{W}_{i,i+1}$to $\ell_{i}$
@@
  • $x_{T}$to $\mu_{j}$
    decision of including $T\in\mathbf{S}$to decision of including $U\in\mathbf{U}^{\prime}$
  • $m^{W,T}_{i,i+1}$to $L_{i}\in U_{j}$
    adjacent singletons in $W$ and $T$to element membership in set
  • $\sum_{W\in\mathbf{W}}c_{W}$to $w_{i}$
@@
Next, comparing the objectives, we can see that Eq. 1 and Eq. 9 have the exact same objective when utilizing the mapping between variables. Finally, we demonstrate a relaxation of Tok’s constraints:
  • Eq. 2 and Eq. 10 are equivalent
@@
Notice that for Tok, we disentangle Eq. 3 and Eq. 4 using the specification of $m^{W,T}_{i,i+1}$ to enable LABEL:{eq:merge-at-most-once}, limiting the covering of an element to one selected set.
Implication.
From our relaxation, it is clear that both Tok and WMC share the same objective while the latter has less constraints and thus can achieve a higher objective value. Knowing that GreedWMC is an $(1-1/e)$-approximate solution to WMC tells us that it would also achieve at least $(1-1/e)$ the maximum attainable objective value on the same problems in Tok. Note that while it is possible that GreedWMC attains an even higher value than the optimum for Tok, because constraints were dropped, the solution obtained by GreedWMC may not correspond to a valid tokenization. Despite this, if GreedTok achieves an objective value that is close to GreedWMC, then one can infer that GreedTok achieves a similar $(1-1/e)$ approximation ratio for the tokenization problem.
4.3 A polynomial-time greedy algorithm
We now informally describe our GreedTok algorithm in a succinct and intuitive manner.  [^2] See Appendix D for a formal pseudocode of how to produce and use $\mathbf{S}$. There are two parts to this: (1) choosing $\mathbf{S}$ from $\mathbf{T}$, and (2) tokenizing words $\mathbf{W}$ using $\mathbf{S}$.
We begin by constructing the candidate token set $\mathbf{T}$ by considering all substrings of length at least 2 within the words $\mathbf{W}$ in the corpus. Then, for any $\mathbf{S}\subseteq\mathbf{T}$, let us define $f(\mathbf{S})$ as the MIP (see Section 4.2) objective value attained by $\mathbf{S}\subseteq\mathbf{T}$. Starting with $\mathbf{S}=\emptyset$, we simply iterate through all $T\in\mathbf{T}\setminus\mathbf{S}$ to find the token
$\tau=\operatorname*{argmax}_{\begin{subarray}{c}T\in\mathbf{T}\setminus\mathbf{S}\\ \text{subject to MIP constraints}\end{subarray}}f(\mathbf{S}\cup\{T\})-f(\mathbf{S})\;,$
to add $\tau$ to $\mathbf{S}$, until $|\mathbf{S}|=k$. Observe that this greedy process induces a natural token ordering amongst the tokens within $\mathbf{S}$.
A direct implementation yields a runtime of $O(|\mathbf{T}|\cdot k\cdot\sum_{W\in\mathbf{W}}|W|)$ when selecting $\mathbf{S}$ and $O(|W|^{2}\cdot\log|W|)$ when tokenizing a word $W$. Note that this token ordering arises from the greedy nature of GreedTok but is not required for solving Tok, just as merge sequences are not fundamental to tokenization. Despite the higher asymptotic costs than BPE, we show in Section 5.1 that with implementation optimizations, GreedTok is practical for real-world NLP applications.
Comparing to BPE.
GreedTok’s token order resembles the merge sequence in BPE, as both select one token per iteration. However, GreedTok operates without the constraints of pairwise merges, allowing more flexible token selection. BPE beats GreedTok in terms of computational complexity, with a selection cost of $O(k\cdot\sum_{W\in\mathbf{W}}|W|)$ and per-word tokenization cost of $O(|W|^{2})$ when using the pairwise caching approach [Ope23]. However, this selection cost is a one-off cost that does not affect downstream applications. Additionally, we empirically show that the modest overhead of $O(\log|W|)$ in tokenization is worth the improvements in downstream tasks; see Section 5.
Comparing to Unigram.
Unigram’s likelihood objective $\mathcal{L}$ can be interpreted as a negative log-weighted version of Tok; see derivation in Appendix A and example where optimizing $\mathcal{L}$ may yield unfavorable behavior. While both GreedTok and Unigram freely select tokens from $\mathbf{T}$, Unigram prunes $\mathbf{T}$ to size $|\mathbf{S}|=k$, while GreedTok builds $\mathbf{S}$ up from $\emptyset$. Computational complexity wise, Unigram’s selection cost of $O(|\mathbf{T}|\cdot\log k\cdot\sum_{W\in\mathbf{W}}|W|)$ beats GreedTok’s but its per-word tokenization cost of $O(k\cdot|W|)$ exceeds GreedTok’s when $k\gg|W|\log|W|$. With large $k$ being common in practical real-world use cases, this is one reason why Unigram is often not used in production systems despite being known to produce higher-quality tokens [BD20, SRZ+24].
5 Empirical evaluation of GreedTok on real-world datasets
5.1 Evaluating GreedTok’s compression
Dataset $|\mathbf{W}|$ $\sum_{W}^{\mathbf{W}}$ count($W$) $|\mathbf{T}|$ max $|\mathbf{S}|=k$ Runtime
UN 105,505 36,985,645 884,708 5,000 6s
[…]
wiki 8,769,943 2,948,721,002 93,571,695 10,000 11m
PubMed 6,527,614 4,149,387,955 121,163,347 10,000 11m
Table 2: Dataset statistics and the time taken for compute with word counts as inputs, conducted with AMD EPYC 9654 @ 2.40GHz. Refer to Section C.1 for additional dataset descriptions.
We compared GreedTok against BPE and Unigram by measuring compression performance across four real-world corpora at varying token budget levels $|\mathbf{S}|=k$; see Table 2. We define the singleton set $\mathbf{B}$ as all 256 byte values, $\mathbf{W}$ as the set of space-delimited UTF-8 strings (byte sequences) extracted from each corpus, and introduce a special token to mark the start of a string following a space character. The function count maps each $W\in\mathbf{W}$ to its frequency in the corpus in UTF-8 format. The candidate token set $\mathbf{T}$ for GreedTok includes all substrings of words in $\mathbf{W}$ while Unigram’s is at the character level. In contrast, BPE begins with an empty $\mathbf{T}$ and builds tokens incrementally from $\mathbf{B}$, depending on the final $|\mathbf{S}|=k$. Thus, both GreedTok and BPE produce final token sets of size $|\mathbf{B}|+k$. However, in addition, we allow Unigram to include frequent (possibly multibyte) characters in the final vocabulary due to the sentencepiece implementation, i.e. it has a final token set size larger than $|\mathbf{B}|+k$.
$k$ 1000 2000 3000 4000 5000 2000 4000 6000 8000 10000
GTK Tokens/Word UN 1.607 1.374 1.268 1.205 1.163 PubMed 1.603 1.397 1.301 1.244 1.206
GTK’s Improvement (%) 4.86 3.99 3.33 2.92 2.54 2.85 2.38 2.02 1.75 1.52
[…]
GTK’s Improvement (%) 5.12 4.94 4.15 3.41 2.89 2.26 1.98 1.71 1.53 1.37
GTK’s Improvement (%) 6.74 4.92 4.30 3.96 3.88 5.62 4.43 3.84 3.75 3.70
Discussion. We see from Table 3 that GreedTok consistently uses fewer tokens on average to represent the same data. Since BPE relies on repeated applications of merge rules to build large tokens, this suggests that many intermediate tokens created during merging may never be used in the final encoding, effectively wasting vocabulary capacity that could be allocated to more useful tokens. Meanwhile, we know from our example in Appendix A that Unigram can over-prioritize whole words at the expense of informative subword tokens, and the empirical results of Table 3 confirms our suspicion that such suboptimal scenarios are not rare.
5.2 Language modeling results
In Appendix B, we show empirical evidence that the token sets produced by GreedTok more closely resemble Unigram than BPE, suggesting that they may inherit some of Unigram’s favorable token characteristics. To test this hypothesis, we pretrain two 1B-parameter language models differing only in tokenizer choice — GreedTok versus BPE. [^3] We use BPE as the baseline, given its status as the most widely adopted tokenizer for LLMs [KHM+23]. Both models use a vocabulary size of 65,536 and are trained on approximately 20% of the DCLM Dedup dataset [TGD+24, LFS+24], with their final token sets being 75% similar.
Pretraining specifics.
Our pre-training corpus is DCLM full-deduped dataset. [^4] Available at: https://huggingface.co/datasets/Zyphra/dclm-dedup. For model training, we use the Dolomite Engine [Mis24]. Our model architecture is a 40-layer Transformer [VSP+17], with embedding size 1536, MLP using SwiGLU activation [Sha20] with intermediate size of 4096, and GQA [ALTdJ+23] layers with 12 query heads and 4 pairs of KV-heads. We used a fixed context length of 4096 tokens and a batch size of $2^{22}\approx 4\mathrm{M}$ tokens. We train GreedTok and BPE tokenizers on approximately 20% of the dataset, randomly selected. For BPE, we use the popular implementation from Huggingface’s tokenizer API. [^5] Available at: https://github.com/huggingface/tokenizers. During training, we follow the same learning rate schedule as [SSM+24], training on the dataset in two phases: the Power Scheduler [SSM+24] in phase 1 and a learning rate with exponential decay in phase 2. However, our phases 1 and 2 use a similar data mixture, sampling the first 500M documents for phase 1, and the next 100M documents for phase 2, which always use 20% of the training iterations of phase 1. We run our experiments using NVIDIA H100 80GB HBM3 cluster, with 96 logical CPU count, training at a rate of $\sim$400B tokens/day.
Methodology.
We compare the BPE-based model (BPEM) against two versions (GTET and GTEP) of the GreedTok-based model under different training constraints, summarized in Table 4:
  1. 1.
    GreedTok Equal Tokens (GTET). Trained using the same number of tokens as BPEM. This setting isolates the impact of denser token representations by fixing the token count. Evaluating GTET is analogous to a setting where model training is compute-constrained, and users can use more text data for training.
  2. 2.
    GreedTok Equal Proportion (GTEP). Trained using the same proportion of the original dataset as BPEM. Here, the number of training tokens differs, based on each tokenizer’s compression ratio, allowing us to examine the effect of using fewer tokens from equivalent text coverage. Evaluating GTEP is analogous to a setting where model training is data-constrained, and users have a limited amount of text for training.
BPEM BPE $8.94\cdot 10^{11}$ $6.29\cdot 10^{11}$ 70.35%
Equal Tokens (GTET) GreedTok $7.35\cdot 10^{11}$ $6.29\cdot 10^{11}$ 85.58%
Table 4: The token count statistics for all three settings. GreedTok uses nearly 18% fewer tokens to represent the entire DCLM Dedup dataset. With 125,000 + 25,000 total iterations in phase 1 and phase 2, the total training tokens for BPEM and GTET come up to about 629B tokens, while we adjust the number of iterations according to the 18% ratio above to match the amount of text trained for the GTEP setting.
In the BPEM and GTET settings, we train the model for 125,000 and 25,000 iterations in phases 1 and 2 respectively. For GTEP, we take the model checkpoint of GTET at the 100,000th training iteration step, followed by an additional 20,000 training iterations in phase 2. The token counts in Table 4 are the results of this process. Due to the constraints of the framework and checkpoint intervals, our dataset proportion of GTEP is not exactly equal to BPEM, but at 68.47% versus 70.35% and favoring BPEM, these are sufficiently close for comparisons.
Evaluation.
We use the popular Language Model Evaluation Harness [GTA+24] toolkit and their predefined evaluation settings to evaluate BPEM, GTET, and GTEP. Several popular benchmark sets were used for evaluation, refer to Section C.2 for more information.
(a) Comparing along number of tokens trained.
(b) Comparing along amount of text trained.
Figure 2: We plot the bits/byte improvement across phase I training for model using GreedTok and BPE on different scales. Since both GTET and GTEP are equivalent in phase 1 for the first 100,000 steps, we examine bits/byte improvement on Wikitext with different scales on the x-axes.
Accuracy (normalized) Accuracy bits/byte
Hella- LAMB- Wino-
BPEM 36.2 67.9 65.6 40.0 75.7 89.8 65.8 81.0 61.1 36.4 62.8 62.0 0.7066
GTEP 37.6 68.8 64.9 39.6 75.6 90.0 67.6 79.0 63.9 36.8 63.5 62.5 0.7028
GTET 38.3 70.0 65.7 40.6 75.8 90.5 67.7 82.0 64.6 37.7 62.6 63.2 0.6989
Discussion.
Previous works report either similar or a decrease in performance stemming from better compression [Gal19, GCE+24, SRZ+24, AFT+24]. One plausible explanation for these results is that, given any sentence, a tokenizer with a lower compression rate uses more tokens, which results in a higher number of total activations in a transformer during inference. This increase in the effective width of the transformer’s computation circuit can increase its expressive power [PMB24], resulting in better performance for poorer compression. However, the results in Table 5 show that GreedTok, compared to its BPE counterpart, has achieved better compression while still maintaining model performance. This suggests an example of meaningful compression, with both GTET and GTEP outperforming BPEM on the evaluated benchmarks. From Fig. 2, when training on equal token count or on equal amount of text, GreedTok is ahead. Our results suggest that the higher compression rate of GreedTok does not negatively affect downstream performance. Furthermore, it is even possible to achieve the same results with GreedTok while using fewer tokens for training.
Figure 3: Plots showing exact $d_{\text{inst}}$ of each problem instance at different $|\mathbf{S}|=k$. As $k$ increases, the ratio of objectives $d_{\text{inst}}$ between $\textsc{GreedTok}(k)$/$\textsc{BPE}(k)$/$\textsc{Unigram}(k)$ and $\textsc{GreedWMC}(k)$ closes, i.e. $d_{\text{inst}}\rightarrow 1$.
5.3 Towards understanding GreedTok’s approximability
We reformulated Problem 1 into a MIP in Section 4.2 as it relaxes naturally into the maximum coverage problem, which has a corresponding $(1-1/e)$ approximate algorithm GreedWMC. Using GreedWMC on the same problem instances of similar $k$, we can calculate the ratio of objectives between GreedTok and GreedWMC and define $d_{\text{inst}}=\frac{\text{{GreedTok}}}{\text{{GreedWMC}}}$ for each instance. Therefore, GreedTok attains an objective value at least $d_{\text{inst}}(1-1/e)$ times the optimal objective of Eq. 1 by definition, with GreedWMC’s objective value is at least that of Tok; see Section 4.2.1.
Discussion. From Fig. 3, for the four selected corpora, we plot $d_{\text{inst}}$ against $k$, observing that as $k$ increases, $d_{\text{inst}}$ climbs towards 1. In addition, to ensure that the results are generalizable to the wider internet corpus, we evaluate on RefinedWeb corpus [PMH+23], and obtain similar findings; see Fig. 4. Empirically, GreedTok achieves an objective value of at least $0.9(1-1/e)$ of the optimal for large $k$, relevant for practical NLP scenarios. With $d_{\text{inst}}\rightarrow 1$, this implies that the room for possible compression improvements narrows.
Figure 4: We sampled documents from RefinedWeb corpus at a probability of 0.01 across 40 independent runs, with mean $d_{\text{inst}}$ plots trending towards 1, empirically showing GreedTok is a $0.9(1-1/e)$-approx. algo.
6 Conclusion
In this work, we showed that the tokenization problem is NP-hard and provided a greedy algorithm GreedTok that is a practical alternative over incumbents BPE and Unigram, and may even be a better option for language pre-training. Although recent works [HKM+24, GT24] had pushed the limits of context length, plausibly reducing the importance of compression, GreedTok can still offer a flexible platform to explore new alternate objectives, such as integrating NLP downstream objectives [BD20] and fairness [LBG+24] constraints into its MIP formulation. Finally, recall that the tokenization problem has the confounding property of being neither supermodular nor submodular. Although we show that GreedTok achieves an approximation ratio of at least $0.9(1-1/e)$ for large $k$, a formal proof is lacking. Nevertheless, this is an intriguing theoretical problem. We hope that our formulation of the tokenization problem and the accompanying toolkit will be valuable for future research.
Computational feasibility.
While the theoretical runtime of GreedTok for selecting $\mathbf{S}$ is $O(|\mathbf{T}|\cdot k\cdot\sum_{W\in\mathbf{W}}|W|)$, a key optimization is to update a token $T$’s marginal contribution only when it is being evaluated for inclusion. This lazy evaluation strategy reduces the average runtime to $\Theta(|\mathbf{T}|\cdot\sum_{W\in\mathbf{W}}|W|)$, making GreedTok practical for large-scale NLP workloads. Table 2 summarizes the time for GreedTok to compute $\mathbf{S}$ at the largest tested size $|\mathbf{S}|=\max k$ (see Section 5.1). In a larger experiment ($|\mathbf{W}|=14.3$M and $|\mathbf{T}|=251$M), GreedTok computed $\mathbf{S}$ in 34 minutes using 160GB of RAM. This cost can be reduced by reducing the search space, limiting $\max|W|$, filtering $\mathbf{W}$, or pruning $\mathbf{T}$ by substring length or frequency. To benchmark encoding speed, we tokenize a subset of the wiki corpus (70K articles, 97M words) using a vocabulary of $|\mathbf{S}|=100$K (from cl100k_base in tiktoken [Ope23]). Our current implementation of GreedTok achieves 700K–800K words per second per thread, and we expect further optimization is possible. These results demonstrate that GreedTok is feasible for integration into modern NLP pipelines.
Acknowledgements
This research/project is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-PhD-2023-08-055T and AISG-PhD/2021-08-013).
Appendix A Tok’s relation to Unigram

### v3

@@ Appendix A Example illustrating that Problem˜1 is neither submodular nor supermodular
In the context of the tokenization problem, the set $\mathbf{T}$ represents the candidate set of all possible tokens. Unfortunately for us, Problem˜1 is neither submodular nor supermodular; see Table˜5.
Table 5: The above table shows that the objective function of Problem˜1 is neither supermodular nor submodular. Suppose we wish to tokenize the word $W=\text{scaredy}$ with candidate token set $\mathbf{T}=\{\text{care, edy, scar, scared, dy}\}$ and singletons $\{\text{s,c,a,r,e,d,y}\}$, and the function $f$ outputs the smallest possible number of final tokens used to represent $W$, i.e. the objective function of Problem˜1 on a single word corpus. Observe that $\mathbf{X}\subseteq\mathbf{Y}\subseteq\mathbf{T}$, $Z\in\mathbf{T}\setminus\mathbf{Y}$, $\mathbf{X}\subseteq\mathbf{Y}^{\prime}\subseteq\mathbf{T}$, and $Z^{\prime}\in\mathbf{T}\setminus\mathbf{Y}^{\prime}$. In case 1, using $\mathbf{X}$ to tokenize $W$ results in using 4 tokens (s, care, d, y)($\text${s, care, d, y}) and one can check that using $\mathbf{X}\cup\{Z\}$ also results in 4 tokens. On the other hand, using $\mathbf{Y}$ results in 4 tokens (s, care, d, y)($\text${s, care, d, y}) but using $\mathbf{Y}\cup\{Z\}$ results in 2 tokens (scar, edy)($\text${scar, edy}). Therefore, $f(\mathbf{X}\cup\{Z\})-f(\mathbf{X})>f(\mathbf{Y}\cup\{Z\})-f(\mathbf{Y})$ and thus $f$ is not supermodular. On the other hand, in case 2, using $\mathbf{Y}^{\prime}$ and $\mathbf{Y}^{\prime}\cup\{Z^{\prime}\}$ to tokenize $W$ results in 2 tokens (scared, y)($\text${scared, y}) while using $\mathbf{X}\cup\{Z^{\prime}\}$ results in 3 tokens (s, care, dy)($\text${s, care, dy}). Therefore, $f(\mathbf{X}\cup\{Z^{\prime}\})-f(\mathbf{X})<f(\mathbf{Y}^{\prime}\cup\{Z^{\prime}\})-f(\mathbf{Y}^{\prime})$ and thus $f$ is not submodular.
@@
$\mathbf{Y}=\{\text{care, edy}\}$ $f(\mathbf{Y})=4$ $\mathbf{Y}^{\prime}=\{\text{care, scared}\}$ $f(\mathbf{Y}^{\prime})=f(\mathbf{Y}^{\prime}\cup\{Z^{\prime}\})=2$
@@
$0=f(\mathbf{X}\cup\{Z\})-f(\mathbf{X})>f(\mathbf{Y}\cup\{Z\})-f(\mathbf{Y})=-2$ $1=f(\mathbf{X}\cup\{Z^{\prime}\})-f(\mathbf{X})<f(\mathbf{Y}^{\prime}\cup\{Z^{\prime}\})-f(\mathbf{Y}^{\prime})=0$
Appendix B Mixed integer program formulation
In this section, we give full details of our mixed integer program (MIP) formulation and provide examples for better understanding.
To formulate Problem˜1 as an MIP, our goal is to choose a subset $\mathbf{S}\subseteq\mathbf{T}$ of size $|\mathbf{S}|\leq k$ such that the following objective is maximized, encoding $\max\sum_{W\in\mathbf{W}}\textsc{count}(W)\cdot\textsc{cover}(W,\mathbf{S})$:
@@ Appendix B Mixed integer program formulation
We remark that the objective Eq.˜2 can be re-expressed as $\max\sum_{T\in\mathbf{T}}\sum_{W\in\mathbf{W}}\sum_{i=1}^{|W|-1}c_{W}\cdot m^{W,T}_{i,i+1}$, making Eq.˜5 redundant. However, this current formulation is useful for showing how to relax Tok to WMC later.
Now, let us interpret and explain the constraints. Eq.˜3 models the constraint that we are choosing a subset of size $|\mathbf{S}|\leq k$. Eq.˜4 models the constraint that we can only use $T\in\mathbf{T}$ to cover if it is chosen in $\mathbf{S}$. Eq.˜5 models the constraint that if a cover happened between two adjacent singletons, then a relevant $T\in\mathbf{T}$ must have been chosen in $\mathbf{S}$. However, Eq.˜6 models the constraint of only covering two adjacent singletons with a single relevant $T\in\mathbf{S}$. Eq.˜7 models the constraint of covering the entire substring $T\in\mathbf{T}$, or leave it uncovered. Eq.˜8 and Eq.˜9 model the constraints preventing the chosen substring $T\in\mathbf{T}$ from sharing the cover with another partially overlapping $T$.
@@
Consider the word $W=\text{ababa}$ and the token $T=\text{aba}$ has $x_{T}=1$, i.e. $T\in\mathbf{S}\subseteq\mathbf{T}$. If we only use $T$ to cover singletons in $W$ with left-to-right priority, then the resultant tokenized form of $W$ is ababa. So, $m^{W}=(1,1,0,0)$, $m^{W,T}=(1,1,0,0)$, and $m^{W,T^{\prime}}=(0,0,0,0)$ for all $T^{\prime}\in\mathbf{T}\setminus\{T\}$. Observe that the 0 bits in $m^{W}$ precisely denote the partitioning positions within $W$. Furthermore, the constraints Eq.˜6 and Eq.˜7 ensure that $T$ is the only token that occupies the first two adjacent singletons, while constraints Eq.˜8 and Eq.˜9 prevent an invalid overlap of $T$ for the last two adjacent singletons. Now, suppose if we also have $T^{\prime}=\text{ba}$ with $x_{T^{\prime}}=1$. Using both $T$ and $T^{\prime}$ to tokenize $W$ results in ababa with $m^{W}=(1,1,0,1)$, $m^{W,T}=(1,1,0,0)$, $m^{W,T^{\prime}}=(0,0,0,1)$, and $m^{W,T^{\prime\prime}}=(0,0,0,0)$ for all $T^{\prime\prime}\in\mathbf{T}\setminus\{T,T^{\prime}\}$.
@@
Appendix C Relation to the weighted maximum coverage problem
In this section, we provide details on how our MIP formulation in Section˜4.1 naturally relaxes into the well known weighted maximum coverage problem (WMC).
Given a set of unique elements $\mathbf{L}=\{L_{1},\dots,L_{|\mathbf{L}|}\}$ and their corresponding weights $\mathcal{W}=\{w_{1},\dots,w_{|\mathbf{L}|}\}$, a collection of sets $\mathbf{U}=\{U_{1},\dots,U_{|\mathbf{U}|}\}$ where $U\in\mathbf{U}\subseteq\mathbf{L}$, and a number $k$, we want to find a subset $\mathbf{U}^{\prime}\subseteq\mathbf{U}$ such that $|\mathbf{U}^{\prime}|\leq k$ and the total weights of covered elements $\sum_{L_{i}\in\bigcup\mathbf{U}^{\prime}}w_{i}$ is maximized. Formulating WMC as a mixed integer program, we have the objective:
@@ Appendix C Relation to the weighted maximum coverage problem
Let us now interpret and explain the constraints. Eq.˜11 limits the number of selected sets $\leq k$. Eq.˜12 ensures that if an element is covered, at least one of the sets containing the element must be included in $\mathbf{U}^{\prime}$. To see that WMC is a relaxation of Tok, we first establish a mapping between the variables between Tok and WMC:
  • $m^{W}_{i,i+1}$ $\to$ $\ell_{i}$
@@
  • $x_{T}$ $\to$ $\mu_{j}$
    decision of including $T\in\mathbf{S}$ $\to$ decision of including $U\in\mathbf{U}^{\prime}$
  • $m^{W,T}_{i,i+1}$ $\to$ $L_{i}\in U_{j}$
    adjacent singletons in $W$ and $T$ $\to$ element membership in set
  • $\sum_{W\in\mathbf{W}}c_{W}$ $\to$ $w_{i}$
@@
Next, comparing the objectives, we can see that Eq.˜2 and Eq.˜10 have the exact same objective when utilizing the mapping between variables. Finally, we demonstrate a relaxation of Tok’s constraints:
  • Eq.˜3 and Eq.˜11 are equivalent
@@
Notice that for Tok, we disentangle Eq.˜4 and Eq.˜5 using the specification of $m^{W,T}_{i,i+1}$ to enable LABEL:{eq:merge-at-most-once}, limiting the covering of an element to one selected set.
Appendix D Tok’s relation to Unigram